OpenStudy (anonymous):

Use your knowledge of the derivative to compute the limit given below: \lim_{h \rightarrow 0} \frac{(x + h)^{3.75} - x^{3.75}}{h} . The derivative that is being calculated is \frac{dy}{dx} = .

OpenStudy (ash2326):

\[\lim_{h \rightarrow 0} \frac{(x + h)^{3.75} - x^{3.75}}{h} \] This is the derivative of \[x^{(3.75)}\] We know that \[\frac{d}{dx} x^n= nx^{n-1}\] so \[\huge \frac{d}{dx} x^{3.75}= 3.75x^{3.75-1}=3.75^x{2.75}\]

OpenStudy (ash2326):

Sorry the last term is \[\huge 3.75 x^{2.75}\]

OpenStudy (anonymous):

thanks